OpenStudy (anonymous):

can someone show me how to take the partial derivative of tan^-1[(2xy)/(x^2-y^2)] with respect to X and Y. I need to prove that it satisfies the Laplace equation. I understand the partial of X and the partial of y must equal zero. I have been working on the partials for hours and it is not working out. can someone please help!

OpenStudy (anonymous):

good grief what algebra. let me get pencil and paper and see if i can get the partials

OpenStudy (anonymous):

thank you... 4 pages of work, what a mess

OpenStudy (anonymous):

did you get \[d/dx = -\frac{2y}{x^2+y^2}\]

OpenStudy (anonymous):

and other partial is similar \[\frac{2x}{x^2+y^2}\]

OpenStudy (anonymous):

the algebra is a big fat mess

OpenStudy (anonymous):

d/dx=[(-2y(x^2+2y)]/(x^2+y^2)^2 ... not sure if it correct

OpenStudy (zarkon):

Satellite has the correct answers

OpenStudy (anonymous):

hard to write all algebra. you have to use \[\frac{1}{(\frac{2xy}{x^2-y^2})^2+1}\times \frac{\delta}{\delta x}\frac{2xy}{x^2-y^2}\]

OpenStudy (anonymous):

& then do the same for y. Answers should add up to 0.

OpenStudy (anonymous):

second term there is \[-\frac{2y(x^2+y^2)}{(x^2-y^2)^2}\]

OpenStudy (anonymous):

i am sorry i cannot be more help. it is almost impossible to write this out. fairly certain of the answer though, so maybe you can work towards it

OpenStudy (anonymous):

and other partial is similar. first term is the same, and second term is \[\frac{2x(x^2+y^2)}{(x^2-y^2)^2}\]

OpenStudy (anonymous):

brb